Solve the quadratic equation: 4x^2=13x+12
Answer by checkley75(3666)

4X^2=13X+12
4X^2-13X-12=0
(4X+3)(X-4)=0
4X+3=0
4X=-3
X=-3/4 OR -.75 ANSWER.
X-4=0
X=4 ANSWER.